# The curvaton ACTs again

### v3

@@ Abstract
The latest results from the Atacama Cosmology Telescope Collaboration have moved the preferred perturbation spectral index $n_{\rm s}$ towards one. We reanalyse constraints on the simplest version of the curvaton model, using $n_{\rm s}$ and the tensor-to-scalar ratio $r$. We show that in the massless curvaton case the model gives the same locus of $n_{\rm s}$$r$ predictions as the power-law model $V(\phi)\propto\phi^{p}$, but with a different and more elegant physical interpretation. The model gives an excellent account of current observational data, both in the regime where the curvaton dominates the observed perturbations and with an admixture of inflaton-originated perturbations of up to about one-third of the total. Addition of the newest South Pole Telescope SPT-3G D1 dataset maintains this conclusion in favor of the curvaton model. Forthcoming large-scale structure surveys, such as that of the recently-launched SPHEREx probe, could discriminate between curvaton and single-field inflation models using the bispectrum.
@@ I Introduction
The analysis of cosmological constraints by the Atacama Cosmology Telescope (ACT) collaboration, following their DR6 data release, has indicated a significant shift upwards in the preferred values of the scalar spectral index $n_{\rm s}$, from the Planck 2018 value of $n_{\rm s}=0.9651\pm 0.0044$ Aghanim et al. (2020) to $n_{\rm s}=0.9743\pm 0.0034$ Louis et al. (2025). The new central value is just over 2-sigma from the Planck result. While within the realm of plausible statistical fluctuations, this change has substantial implications for inflationary modelling.
@@ II.1 Model parameters and density perturbations
$V(\phi,\sigma)=\frac{1}{2}m_{\phi}^{2}\phi^{2}+\frac{1}{2}m_{\sigma}^{2}\sigma^{2}.$ (1)
The inflaton $\phi$ is responsible for driving inflation via its potential energy, while the curvaton $\sigma$ acquires perturbations during inflation that subsequently convert to adiabatic when the curvaton decays to normal matter after inflation ends. We will closely follow the detailed analysis we made in Ref. Byrnes et al. (2014), which allows an arbitrary admixture of inflaton and curvaton contributions to the observed adiabatic power spectrum. A similar recent analysis looking at other choices of the inflaton potential was made in Ref. Choi et al. (2025).
@@
in the single-field model. Taking the observed amplitude as Aghanim et al. (2020)
@@
where henceforth $N_{*}$ is the $e$-foldings number at which the Planck normalization scale $0.05\,{\rm Mpc}^{-1}$ crosses the Hubble radius during inflation. The required $r_{\rm dec}$ value fixes the decay timescale of the curvaton. Evaluating Eq. (10) requires knowledge of the full curvaton evolution, but in practice we will only use $r_{\rm dec}$ via Eq. (12) as a constraint on model parameters by requiring that it takes the physically realisable values $10<r_{\rm dec}<$ 1. This also allows a simple generalisation to models where the inflaton decays partly to curvatons which later themselves decay into the thermal bath Byrnes et al. (2016). We will see that the lower bound is tightened by the constraint on local $f_{\rm NL}$.
@@ II.2 Parametrization of the number of $e$-foldings
To impose accurate constraints we need to identify the correct number of $e$-foldings corresponding to the $k_{*}=0.05\,{\rm Mpc}^{-1}$ pivot scale at which observables are evaluated. This is given by Liddle and Leach (2003); Byrnes et al. (2014)
$N_{\rm*}=58+\frac{1}{4}\ln\epsilon+\frac{1}{4}\ln\frac{V_{\rm hor}}{\rho_{\rm end}}+\frac{1}{12}\ln\frac{\rho_{\rm reh}}{\rho_{\rm end}}\,,$ (13)
where all quantities are as in single-field models. We also parametrize the total amount of reheating $e$-foldings, given by the last term of Eq. (13), as $N_{\rm matter}$ which includes the reheating of both the inflaton and the curvaton, obtaining
@@
We note that this relation sets a firm upper limit of $58$ as the number of $e$-foldings corresponding to the Planck pivot scale.
@@
Figure 1: Observational constraints in the $n_{\rm s}$$r$ plane from Ref. Calabrese et al. (2025). The combined Planck, ACT and BICEP/Keck CMB spectral data, enhanced by CMB lensing and by baryon acoustic oscillation data from DESI DR1, are shown in purple. The three dashed lines show the power-law potentials for $e$-folding choices $N_{*}=40$, $50$, and $60$, which are also the predictions of our curvaton model in the massless curvaton limit. The sets of symbols, from top to bottom, correspond to $f=1/2$, $1/3$, and $1/6$. [Adapted from Fig. 10 of Ref. Calabrese et al. (2025), under Creative Commons BY 4.0 License.]
@@ II.3 Observables: Linear power spectrum
If there is a single period of inflation, $\epsilon\simeq\eta_{\phi}\simeq 1/2N_{*}$. [^3] This corrects a typo in Ref. Byrnes et al. (2014) which fortunately was not propagated to any of the results.
@@
This expression allows for ready distinction between the contribution of the inflaton versus curvaton for each model spectrum, as a function of the inflaton contribution $f$. It reduces to each of these two regimes for $f\sim 1$ and $f\ll 1$ respectively. We further note that $n_{\rm S}$ has no dependence on $\sigma_{*}$. Lastly, Eq. (17) has an implicit dependence on the number of $e$-foldings and the chosen pivot scale, which we presented in Section II.2.
@@
The standard reference guide in discussing observational constraints on inflation is the power-law potential $V(\phi)\propto\phi^{p}$, whose predictions in the $n_{\rm s}$$r$ plane are seen in Fig. 1 for several $N_{*}$ values. While early inflation modelling envisaged an even integer power such as $p=2$ or $4$, observational constraints, particularly the non-detection of $r$, have now driven the limit on $p$ well below one, making the justification and interpretation of the potential much harder. But remarkably, in the limit of negligible curvaton mass, the curvaton model predicts exactly the same locus in the $n_{\rm s}$$r$ plane, as we now show.
@@
This lets the curvaton model match any $p$ in the range 00 to $2$, passing between curvaton and inflaton domination of the generated perturbations respectively. In particular, the curvaton model can exactly mimic the combined $(n_{\rm s},r)$ values of $p<1$ while using only simple quadratic potentials.
@@
Despite the match in the $n_{\rm s}$$r$ predictions, the running of the spectral index will not be identical, as a change in $N_{*}$ changes the value of $m_{\rm single}$ and hence $f$ [^4] We thank David Wands for pointing this out to us. The runnings $\alpha\equiv dn_{\rm s}/d\ln k$ for the $\phi^{p}$ model Kosowsky and Turner (1995); Chiba and Kohri (2014) and the curvaton model Byrnes and Wands (2006) are given, respectively, by
@@ II.4 Observables: Non-gaussianity
Taking the Planck 95% confidence observational upper limit of $f_{\rm NL}<9.3$ Akrami et al. (2020) leads to $r_{\rm dec}>0.11$ in the curvaton limit and $r_{\rm dec}>0.05$ if the inflaton contributes 33% of the perturbations. We note that values of $r_{\rm dec}$ quite close to but not equal to one require rather fine-tuned parameters, and hence $r_{\rm dec}\simeq 1$ should arguably be considered the natural expectation Hardwick and Byrnes (2015); Torrado et al. (2018). We therefore expect $f_{\rm NL}=-5/4$ if the curvaton dominates, a value which SPHEREx, which launched in March 2025, may be able to distinguish from gaussianity in the next few years Doré et al. (2014); Heinrich et al. (2024). A detection of $f_{\rm NL}<-5/4$ would completely rule out the quadratic curvaton scenario independently of the inflaton potential, whilst a constraint of e.g. $|f_{\rm NL}|<0.5$ would strongly disfavour the curvaton scenario over single-field scenarios which gives rise to $f_{\rm NL}\simeq 0$. A detection of $f_{\rm NL}>0$ would rule out single-field inflation and put tight constraints on $r_{\rm dec}$. For example if SPHEREx detected $42<f_{\rm NL}<$ 4 and if constraints on $r$ also significantly strengthened, then this would require $0.320.21<r_{\rm dec}<$ 0.32 for the quadratic curvaton scenario to remain viable.
@@ III Conclusions
The significant shift in $n_{\rm s}$ prompted by the ACT DR6 results Calabrese et al. (2025); Louis et al. (2025) has restored the viability of simple curvaton models. We have uncovered an exact correspondence between the observational predictions of the simplest curvaton model Bartolo and Liddle (2002) and the single-field $\phi^{p}$ potential that is widely used as a benchmark, with the power-law $p$ replaced by the admixture $f$ of inflaton and curvaton contributions to the perturbations. Moreover, the curvaton model uses only simple quadratic potentials, yet can give predictions that mimic $\phi^{p}$ models with $p\ll 1$ that can be otherwise hard to motivate.
@@ Acknowledgements
We thank Renata Kallosh and Andrei Linde for comments, and David Wands for important insights concerning the spectral index running. CB is supported by STFC grants ST/X001040/1 and ST/X000796/1. M.C and A.R.L. were supported by the Fundação para a Ciência e a Tecnologia (FCT) through the research grants UIDB/04434/2020 and UIDP/04434/2020. M.C. acknowledges support from the FCT through the Investigador FCT Contract No. CEECIND/02581/2018 and POPH/FSE (EC), and A.R.L. through the Investigador FCT Contract No. CEECIND/02854/2017 and POPH/FSE (EC). This article/publication is based upon work from COST Action CA21136 – “Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse)”, supported by COST (European Cooperation in Science and Technology).

### v4

# Curvaton in light of the ACT results

@@ Abstract
The latest results from the Atacama Cosmology Telescope Collaboration have moved the preferred perturbation spectral index $n_{\rm s}$ closer towards one. We reanalyse constraints on the simplest version of the curvaton model, using $n_{\rm s}$ and the tensor-to-scalar ratio $r$. We show that in the massless curvaton case the model gives the same locus of $n_{\rm s}$$r$ predictions as the power-law model $V(\phi)\propto\phi^{p}$, but with a different and more elegant physical interpretation. The model gives an excellent account of current observational data, both in the regime where the curvaton dominates the observed perturbations and with an admixture of inflaton-originated perturbations of up to about one-third of the total. Addition of the newest South Pole Telescope SPT-3G D1 dataset maintains this conclusion in favor of the curvaton model. Forthcoming large-scale structure surveys, such as that of the recently-launched SPHEREx probe, could discriminate between curvaton and single-field inflation models using the bispectrum.
@@ I Introduction
The analysis of cosmological constraints by the Atacama Cosmology Telescope (ACT) collaboration, following their DR6 data release, has indicated a consequential shift upwards in the preferred values of the scalar spectral index $n_{\rm s}$, from the Planck 2018 value of $n_{\rm s}=0.9651\pm 0.0044$ Aghanim et al. (2020) to $n_{\rm s}=0.9743\pm 0.0034$ Louis et al. (2025) (in a compilation using DESI-DR1 data Adame et al. (2025a, b, c)). The new central value is just over 2-sigma from the Planck result. While within the realm of plausible statistical fluctuations, this change has substantial implications for inflationary modelling. The higher value has been reinforced in subsequent analyses, for example the revised (v2) ACT paper Louis et al. (2025) includes an analysis updating to DESI-DR2 Abdul Karim et al. (2025) that yields the somewhat higher again $n_{\rm s}=0.9752\pm 0.0030$, while the Dark Energy Survey Y6 release finds a yet higher value $n_{\rm s}=0.977\pm 0.003$ from an all-data compilation including Planck and ACT Abbott et al. (2026).
@@ II.1 Model parameters and density perturbations
$V(\phi,\sigma)=\frac{1}{2}m_{\phi}^{2}\phi^{2}+\frac{1}{2}m_{\sigma}^{2}\sigma^{2}\,$, (1)
and a flat field-space metric. The inflaton $\phi$ is responsible for driving inflation via its potential energy, while the curvaton $\sigma$ acquires perturbations during inflation that subsequently convert to adiabatic when the curvaton decays to normal matter after inflation ends. We will closely follow the detailed analysis we made in Ref. Byrnes et al. (2014), which allows an arbitrary admixture of inflaton and curvaton contributions to the observed adiabatic power spectrum. A similar recent analysis looking at other choices of the inflaton potential was made in Ref. Choi et al. (2025).
@@
in the single-field model. Taking the observed amplitude as Louis et al. (2025)
@@
where henceforth $N_{*}$ is the $e$-foldings number at which the Planck normalization scale $0.05\,{\rm Mpc}^{-1}$ crosses the Hubble radius during inflation. The required $r_{\rm dec}$ value fixes the decay timescale of the curvaton. Evaluating Eq. (10) requires knowledge of the full curvaton evolution, but in practice we will only use $r_{\rm dec}$ via Eq. (12) as a constraint on model parameters by requiring that it takes the physically realisable values $0<r_{\rm dec}<1$. This also allows a simple generalisation to models where the inflaton decays partly to curvatons which later themselves decay into the thermal bath Byrnes et al. (2016). We will see that the lower bound is tightened by the constraint on local $f_{\rm NL}$.
@@ II.2 Parametrization of the number of $e$-foldings
To impose accurate constraints we need to identify the correct number of $e$-foldings corresponding to the $k_{*}=0.05\,{\rm Mpc}^{-1}$ pivot scale at which observables are evaluated. With $V_{\rm hor}$ the potential when the pivot scale crossed the horizon during inflation, $\rho_{\rm end}$ the energy density at the end of inflation, and $\rho_{\rm reh}$ the energy density at the end of reheating, this is given by Liddle and Leach (2003); Byrnes et al. (2014)
$N_{\rm*}=58+\frac{1}{4}\ln(rf)+\frac{1}{4}\ln\frac{V_{\rm hor}}{\rho_{\rm end}}+\frac{1}{12}\ln\frac{\rho_{\rm reh}}{\rho_{\rm end}}\,$, (13)
where all quantities are as in single-field models. The factor $f$ is needed because it modifies the normalization of $V_{\rm hor}$ relative to the observed power spectrum amplitude when the curvaton perturbations are non-negligible [^3] This factor was omitted in Ref. Byrnes et al. (2014), but the error did not propagate., but where we need to include the epoch of curvaton reheating in addition to the usual inflaton reheating. We parametrize the total amount of reheating $e$-foldings, given by the last term of Eq. (13), as $N_{\rm matter}$ which includes the reheating of both the inflaton and the curvaton, obtaining
@@
We note that this relation sets a firm upper limit of $58$ as the number of $e$-foldings corresponding to the Planck pivot scale (see also Ref. Germán et al. (2023)).
@@
Figure 1: Observational constraints in the $n_{\rm s}$$r$ plane from Ref. Calabrese et al. (2025). The combined Planck, ACT and BICEP/Keck CMB spectral data, enhanced by CMB lensing and by baryon acoustic oscillation data from DESI-DR1, are shown in purple (see also Figure 59 in v2 of their article for an update to DESI-DR2). The three dashed lines show the power-law potentials for $e$-folding choices $N_{*}=40$, $50$, and $60$, which are also the predictions of our curvaton model in the massless curvaton limit. The sets of symbols, from top to bottom, correspond to $f=1/2$, $1/3$, and $1/6$. [Adapted from Fig. 10 of Ref. Calabrese et al. (2025), under Creative Commons BY 4.0 License.]
@@ II.3 Observables: Linear power spectrum
If there is a single period of inflation, $\epsilon\simeq\eta_{\phi}\simeq 1/2N_{*}$. [^4] This corrects a typo in Ref. Byrnes et al. (2014), which fortunately was not propagated to any of the results.
@@
This expression allows for ready distinction between the contribution of the inflaton versus curvaton for each model spectrum, as a function of the inflaton contribution $f$. It reduces to each of these two regimes for $f\sim 1$ and $f\ll 1$ respectively. We further note that $n_{\rm S}$ has no dependence on $\sigma_{*}$ at leading order in slow-roll. Lastly, Eq. (17) has an implicit dependence on the number of $e$-foldings and the chosen pivot scale, which we presented in Section II.2.
@@
The standard reference guide in discussing observational constraints on inflation is the power-law potential $V(\phi)\propto\phi^{p}$, whose predictions in the $n_{\rm s}$$r$ plane are seen in Fig. 1 for several $N_{*}$ values. While early inflation modelling envisaged an even integer power such as $p=2$ or $4$, observational constraints, particularly the non-detection of $r$, have now driven the limit on $p$ well below one, making the justification and interpretation of the potential much harder. Even the most-preferred monodromy value of $p=2/3$ Silverstein and Westphal (2008); McAllister et al. (2014) is under strong pressure, especially for lower $N_{*}$ values. But remarkably, in the limit of negligible curvaton mass, the curvaton model predicts exactly the same locus in the $n_{\rm s}$$r$ plane, as we now show.
@@
This lets the curvaton model match any $p$ in the range 0 to $2$, passing between curvaton and inflaton domination of the generated perturbations respectively. In particular, the curvaton model can exactly mimic the combined $(n_{\rm s},r)$ values of $p<1$ while using only simple quadratic potentials.
@@
Despite the match in the $n_{\rm s}$$r$ predictions, the running of the spectral index will not be identical, as a change in $N_{*}$ changes the value of $m_{\rm single}$ and hence ff. [^5] We thank David Wands for pointing this out to us. The runnings $\alpha\equiv dn_{\rm s}/d\ln k$ for the $\phi^{p}$ model Kosowsky and Turner (1995); Chiba and Kohri (2014) and the curvaton model Byrnes and Wands (2006) are given, respectively, by
@@ II.4 Observables: Non-Gaussianity
Taking the Planck 95% confidence observational upper limit of $f_{\rm NL}<9.3$ Akrami et al. (2020) leads to $r_{\rm dec}>0.11$ in the curvaton limit and $r_{\rm dec}>0.05$ if the inflaton contributes 33% of the perturbations. We note that values of $r_{\rm dec}$ quite close to but not equal to one require rather fine-tuned parameters, and hence $r_{\rm dec}\simeq 1$ should arguably be considered the natural expectation Hardwick and Byrnes (2015); Torrado et al. (2018). We therefore expect $f_{\rm NL}=-5/4$ if the curvaton dominates, a value which SPHEREx, which launched in March 2025, may be able to distinguish from Gaussianity in the next few years Doré et al. (2014); Heinrich et al. (2024). A detection of $f_{\rm NL}<-5/4$ would completely rule out the quadratic curvaton scenario independently of the inflaton potential, whilst a constraint of e.g. $|f_{\rm NL}|<0.5$ would strongly disfavour the curvaton scenario over single-field scenarios which gives rise to $f_{\rm NL}\simeq 0$. A detection of $f_{\rm NL}>0$ would rule out single-field inflation and put tight constraints on $r_{\rm dec}$. For example if SPHEREx detected $2<f_{\rm NL}<4$ and if constraints on $r$ also significantly strengthened, then this would require $0.21<r_{\rm dec}<0.32$ for the quadratic curvaton scenario to remain viable.
@@ III Conclusions
The significant shift in $n_{\rm s}$ prompted by the ACT DR6 results Louis et al. (2025) has restored the viability of simple curvaton models. We have uncovered an exact correspondence between the observational predictions of the simplest curvaton model Bartolo and Liddle (2002) and the single-field $\phi^{p}$ potential that is widely used as a benchmark, with the power-law $p$ replaced by the admixture $f$ of inflaton and curvaton contributions to the perturbations. Moreover, the curvaton model uses only simple quadratic potentials, yet can give predictions that mimic $\phi^{p}$ models with $p\ll 1$ that can be otherwise hard to motivate.
@@ Acknowledgements
We thank Renata Kallosh and Andrei Linde for comments. Also, we thank Charlotte Wood for pointing out an error in the original version of Eq. (13), and David Wands for important insights concerning the spectral index running. CB is supported by STFC grants ST/X001040/1 and ST/X000796/1. M.C and A.R.L. were supported by the Fundação para a Ciência e a Tecnologia (FCT) through the research grants UIDB/04434/2020 and UIDP/04434/2020. M.C. acknowledges support from the FCT through the Investigador FCT Contract No. CEECIND/02581/2018 and POPH/FSE (EC), and A.R.L. through the Investigador FCT Contract No. CEECIND/02854/2017 and POPH/FSE (EC). This article/publication is based upon work from COST Action CA21136 – “Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse)”, supported by COST (European Cooperation in Science and Technology).